OpenStudy (anonymous):

find an equation of the line with the giving slope that passes through the giving point. write the equation in the form Ax+By=C m=1/8; (-9,1)

OpenStudy (anonymous):

y - 1 = 1/8 (x - -9) 8y - 8 = x + 9 x + 8y = -17

OpenStudy (anonymous):

sorry its x - 8y = -17

OpenStudy (anonymous):

thank you, I just dont get some of this... I have a few more if thats okay

OpenStudy (anonymous):

check by plugging in the values : -9 - 8(1) = -17

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

find an equation of the line through the pair of points.Write the equation in the form Ax+By=C (-6,3) and (4,0)